# Locating Superspreaders in SI Cascades

### 05 Mar 2024 v1

@@ 4 Cascade dynamics
Let $v\in V$ and suppose that $\deg(v)\lambda^{2}\delta^{2}\leq 1$. Fix $t\geq 0$, and let $v\in V\setminus{\mathcal{I}}[0,t]$. Then
($\operatorname{\mathbb{P}}\left(\left.\{T(v)\in[t,t+\delta]\}\cap{\mathcal{E}}\right|{\mathcal{F}}_{t}\right)\leq 3d\log^{2}(n)\sqrt{\deg(v)\lambda^{2}\delta^{2}}.$
[update ${\mathcal{E}}_{t}$ to ${\mathcal{A}}_{t}$]
@@
${\mathcal{E}}_{t}:=\left\{\text{For all $v\in V$ such that $\deg(v)\geq D$ and$T(v)>t$, it holds that $|{\mathcal{N}}(v)\cap{\mathcal{I}}[0,t]|\leq 2d\sqrt{\deg(v)}\log^{2}n$}\right\}.$
Notice that ${\mathcal{E}}_{t}$ is ${\mathcal{F}}_{t}$-measurable, and ${\mathcal{A}}\subset{\mathcal{E}}_{t}$. In particular,
($\operatorname{\mathbb{P}}\left(\{T(v)\leq t+\delta\}\cap{\mathcal{E}}|{\mathcal{F}}_{t}\right)\leq\operatorname{\mathbb{P}}\left(\{T(v)\leq t+\delta\}\cap{\mathcal{E}}_{t}|{\mathcal{F}}_{t}\right)=\operatorname{\mathbb{P}}(T(v)\leq t+\delta|{\mathcal{F}}_{t})\mathbf{1}({\mathcal{E}}_{t}).$ (4.4)
We will therefore proceed by bounding $\operatorname{\mathbb{P}}(T(v)\leq t+\delta|{\mathcal{F}}_{t})$ under the assumption that ${\mathcal{E}}_{t}$ holds.
Suppose that $\deg(v)\leq d$. Then by Lemma 4.4, $T(v)-t$ is stochastically lower bounded by a $\mathrm{Exp}(\lambda d)$ random variable, so that
@@
Next, suppose that $\deg(v)\geq D$, and let ${\mathcal{P}}_{2}(v)$ be the set of paths of length 2 starting from $v$. Then
@@
$\geq\min\left\{\min_{u\in{\mathcal{N}}(v)\cap{\mathcal{I}}[0,t]}F(u,v),\min_{P\in{\mathcal{P}}_{2}(v)}\mathsf{weight}(P)\right\},$ (4.7)
which can be explained as follows. If the minimizing path $P\in{\mathcal{P}}(v,t)$ in (4.6) has length 1, then the minimum is achieved by $F(u,v)$ for some $u\in{\mathcal{N}}(v)\cap{\mathcal{I}}[0,t]$. Otherwise, the minimizing path in (4.6) has length at least 2, which must include a path of length 2 starting at $v$. In light of (4.7), we see that
$\operatorname{\mathbb{P}}(T(v)\leq t+\delta|{\mathcal{F}}_{t})$ $\leq\operatorname{\mathbb{P}}\left(\left.\min\left\{\min_{u\in{\mathcal{N}}(v)\cap{\mathcal{I}}[0,t]}F(u,v),\min_{P\in{\mathcal{P}}_{2}(v)}\mathsf{weight}(P)\right\}\leq\delta\right|{\mathcal{F}}_{t}\right)$
$\leq\sum_{u\in{\mathcal{N}}(v)\cap{\mathcal{I}}[0,t]}\operatorname{\mathbb{P}}(F(u,v)\leq\delta|{\mathcal{F}}_{t})+\sum_{P\in{\mathcal{P}}_{2}(v)}\operatorname{\mathbb{P}}(\mathsf{weight}(P)\leq\delta|{\mathcal{F}}_{t})$
$\stackrel{{\scriptstyle(a)}}{{\leq}}|{\mathcal{N}}(v)\cap{\mathcal{I}}[0,t]|\lambda\delta+|{\mathcal{P}}_{2}(v)|\lambda^{2}\delta^{2}$
$\stackrel{{\scriptstyle(b)}}{{\leq}}2d\log^{2}(n)\sqrt{\deg(v)\lambda^{2}\delta^{2}}+d\left(\deg(v)\lambda^{2}\delta^{2}\right)$
$\stackrel{{\scriptstyle(c)}}{{\leq}}3d\log^{2}(n)\sqrt{\deg(v)\lambda^{2}\delta^{2}},$ (4.8)
where the inequality $(a)$ follows since $F(u,v)\sim\mathrm{Exp}(\lambda)$ and $\mathsf{weight}(P)\sim\mathrm{Gamma}(2,\lambda)$ for $P\in{\mathcal{P}}_{2}(v)$. The inequality $(b)$ uses the bound $|{\mathcal{N}}(v)\cap{\mathcal{I}}[0,t]|\leq 2d\sqrt{\deg(v)}\log^{2}n$ which holds on the event ${\mathcal{E}}_{t}$, and also uses that the number of paths of length 2 starting from $v$ is at most $d\cdot\deg(v)$, since we assume that high-degree vertices are not neighbors in $G$. Finally, $(c)$ holds by our assumption that $\deg(v)\lambda^{2}\delta^{2}\leq 1$. ∎
Lemma 4.7 can be readily extended to the case where $t$ is replaced by a stopping time. The following result provides one useful example of this idea.
@@
Let $u,v\in V$ and suppose that $\max\{\deg(u),\deg(v)\}\lambda^{2}\delta^{2}\leq 1$. Then
@@
($\operatorname{\mathbb{P}}\left(\{|T(v)-T(u)|\leq\delta\}\cap{\mathcal{E}}\right)\leq 3d\log^{2}(n)\sqrt{\deg(v)\lambda^{2}\delta^{2}}.$
@@
Define the stopping time $T:=\min\{T(u),T(v)\}$. Let $t\geq 0$, and condition on ${\mathcal{F}}_{t}$. Assuming that the ${\mathcal{F}}_{t}$-measurable event $\{T=t\}$ holds and, without loss of generality that $T=T(u)$, we have by Lemma 4.7 that
($\operatorname{\mathbb{P}}\left(\{T(v)\leq t+\delta\}\cap{\mathcal{E}}|{\mathcal{F}}_{t}\right)\mathbf{1}(T=t)\leq\left(3d\log^{2}(n)\sqrt{\deg(v)\lambda^{2}\delta^{2}}\right)\mathbf{1}(T=t).$
Since $T<\infty$ almost surely, we have that
($\operatorname{\mathbb{P}}\left(\{T(v)\leq t+\delta\}\cap{\mathcal{E}}\right)$ $=\operatorname{\mathbb{E}}\left[\operatorname{\mathbb{P}}\left(\{T(v)\leq t+\delta\}\cap{\mathcal{E}}|{\mathcal{F}}_{T}\right)\right]$
[…]
$\leq 3d\log^{2}(n)\sqrt{\deg(v)\lambda^{2}\delta^{2}}.$
5 Second derivative method for $\alpha\in(2/3,1]$
5.1 First derivative is approximately the size of the cut
Conditioned on ${\mathcal{F}}_{t}$, we partition ${\mathcal{I}}[t,t+\delta]={\mathcal{I}}_{1}\cup{\mathcal{I}}_{2}$, where
${\mathcal{I}}_{1}$ :=$:=\left\{v\in\partial{\mathcal{I}}(t):\min_{u\in{\mathcal{N}}(v)\cap{\mathcal{I}}[0,t]}F(u,v)\leq\delta\right\}$
${\mathcal{I}}_{2}$ :=$:=\left\{v\in{\mathcal{I}}[t,t+\delta]:P\in{\mathcal{P}}(v,t)\text{ and }\mathsf{weight}(P)\leq\delta\Rightarrow|P|\geq 2\right\}={\mathcal{I}}[t,t+\delta]\setminus{\mathcal{I}}_{1}.$
We proceed by studying the sizes of ${\mathcal{I}}_{1}$ and ${\mathcal{I}}_{2}$ separately. Our first result characterizes $|{\mathcal{I}}_{1}|$, essentially through an application of Bernstein’s inequality.
@@ 5.1 First derivative is approximately the size of the cut
Let $G$ be a tree on $n$ vertices, let $\beta>0$ and assume that $\gamma>\max\{(1-\beta)/2,1-2\beta\}$. Then for any $t\geq 0$, it holds that
$\operatorname{\mathbb{P}}\left(\left.\left||{\mathcal{I}}_{1}|-\frac{\lambda}{n^{\beta}}|\operatorname{\mathsf{cut}}({\mathcal{I}}[0,t])|\right|\geq n^{\gamma}\right|{\mathcal{F}}_{t}\right)=o\left(\frac{1}{n}\right).$

### 06 Mar 2024 v2

@@ 3 Preliminaries: Cascade dynamics and infection times
Let $v\in V$ and let $\delta>0$ satisfy $\lambda^{2}\delta^{2}n\leq 1$. Fix $t\geq 0$, and let $v\in V\setminus{\mathcal{I}}(t)$. Then
($\operatorname{\mathbb{P}}\left(\left.\{T(v)\in[t,t+\delta]\}\cap{\mathcal{A}}\right|{\mathcal{F}}_{t}\right)\leq 3d\lambda\delta\sqrt{n}\log^{2}(n).$
@@
${\mathcal{A}}_{t}:=\left\{\text{For all $v\in V$ such that $T(v)>t$, it holds that $|{\mathcal{N}}(v)\cap{\mathcal{I}}(t)|\leq 2d\sqrt{n}\log^{2}n$}\right\}.$
Notice that ${\mathcal{A}}_{t}$ is ${\mathcal{F}}_{t}$-measurable, and ${\mathcal{A}}\subset{\mathcal{A}}_{t}$. Moreover, for $v$ such that $T(v)>t$,
($\operatorname{\mathbb{P}}\left(\{T(v)\leq t+\delta\}\cap{\mathcal{A}}|{\mathcal{F}}_{t}\right)\leq\operatorname{\mathbb{P}}\left(\{T(v)\leq t+\delta\}\cap{\mathcal{A}}_{t}|{\mathcal{F}}_{t}\right)=\operatorname{\mathbb{P}}(T(v)\leq t+\delta|{\mathcal{F}}_{t})\mathbf{1}({\mathcal{A}}_{t}).$ (3.2)
We will therefore proceed by bounding $\operatorname{\mathbb{P}}(T(v)\leq t+\delta|{\mathcal{F}}_{t})$ under the assumption that ${\mathcal{A}}_{t}$ holds.
Suppose that $\deg(v)\leq d$. Then by Lemma 3.3, $T(v)-t$ stochastically dominates a $\mathrm{Exp}(\lambda d)$ random variable, so that
@@
Next, suppose that $\deg(v)\geq D$, and let ${\mathcal{P}}_{2}(v)$ be the set of paths of length 2 starting from $v$. Then by Proposition 3.1,
@@
$\geq\min\left\{\min_{u\in{\mathcal{N}}(v)\cap{\mathcal{I}}(t)}F(u,v),\min_{P\in{\mathcal{P}}_{2}(v)}\mathsf{weight}(P)\right\},$ (3.5)
which can be explained as follows. If the minimizing path $P\in{\mathcal{P}}(v,t)$ in (3.4) has length 1, then the minimum is achieved by $F(u,v)$ for some $u\in{\mathcal{N}}(v)\cap{\mathcal{I}}(t)$. Otherwise, the minimizing path in (3.4) has length at least 2, which must include a path of length 2 starting at $v$. In light of (3.5),
$\operatorname{\mathbb{P}}(T(v)\leq t+\delta|{\mathcal{F}}_{t})$ $\leq\operatorname{\mathbb{P}}\left(\left.\min\left\{\min_{u\in{\mathcal{N}}(v)\cap{\mathcal{I}}(t)}F(u,v),\min_{P\in{\mathcal{P}}_{2}(v)}\mathsf{weight}(P)\right\}\leq\delta\right|{\mathcal{F}}_{t}\right)$
$\leq\sum_{u\in{\mathcal{N}}(v)\cap{\mathcal{I}}(t)}\operatorname{\mathbb{P}}(F(u,v)\leq\delta|{\mathcal{F}}_{t})+\sum_{P\in{\mathcal{P}}_{2}(v)}\operatorname{\mathbb{P}}(\mathsf{weight}(P)\leq\delta|{\mathcal{F}}_{t})$
$\stackrel{{\scriptstyle(a)}}{{\leq}}|{\mathcal{N}}(v)\cap{\mathcal{I}}(t)|\lambda\delta+|{\mathcal{P}}_{2}(v)|\lambda^{2}\delta^{2}$
$\stackrel{{\scriptstyle(b)}}{{\leq}}2d\lambda\delta\sqrt{n}\log^{2}(n)+d\lambda^{2}\delta^{2}n$
$\stackrel{{\scriptstyle(c)}}{{\leq}}3d\lambda\delta\sqrt{n}\log^{2}(n),$ (3.6)
where the inequality $(a)$ follows since $F(u,v)\sim\mathrm{Exp}(\lambda)$ and $\mathsf{weight}(P)\sim\mathrm{Gamma}(2,\lambda)$ for $P\in{\mathcal{P}}_{2}(v)$. The inequality $(b)$ uses the bound $|{\mathcal{N}}(v)\cap{\mathcal{I}}(t)|\leq 2d\sqrt{n}\log^{2}n$ which holds on the event ${\mathcal{A}}_{t}$, and also uses that the number of paths of length 2 starting from $v$ is at most $dn$, since we assume that high-degree vertices are not neighbors in $G$ (see Definition 1.1) and the degree of any vertex is at most $n$. Finally, $(c)$ holds by our assumption that $\lambda^{2}\delta^{2}n\leq 1$. ∎
4 Analysis of Algorithm 3: Proof of Theorem 2.1
The main result behind the performance guarantees of Algorithm 3 is the following lemma, which shows that $\widehat{\deg}_{\delta,i}(v)$ is a good approximation for $\lambda\deg(v)$ in at least half of the cascades, for all $v\in V$.
@@
Suppose that $\beta,\gamma$ satisfies $\beta>1/2$ and $\gamma>(1+\beta)/2$. Additionally assume
$K\geq\frac{15}{\beta-1/2}.$
Then with probability $1-o(1)$, for each $v\in V$ there exists a set $S_{v}\subset[K]$ such that $|S_{v}|\geq\frac{K}{2}$ and
$\left|\widehat{\deg}_{\delta,i}(v)-\lambda\deg(v)\right|\leq n^{\gamma},\qquad\forall i\in S_{v}.$
Before proving the lemma, we show how it can be used to prove Theorem 2.1.
Proof of Theorem 2.1.
Let $\gamma\in(\alpha/2+3/8,\alpha)$. It is readily verified that the choice of $\alpha,\beta,\gamma,K$ satisfies the assumptions of Lemma 4.1. In the case where $v\in\mathsf{highdeg}(G)$, it holds that for all $i\in S_{v}$,
$\widehat{\deg}_{\delta,i}(v)\geq\lambda\deg(v)-n^{\gamma}\geq\lambda n^{\alpha}-n^{\gamma}\geq\frac{n^{\alpha}}{\log n}.$
In particular, $\widehat{\deg}_{\delta,i}(v)\geq n^{\alpha}/\log(n)$ for at least half of the cascades. On the other hand, if $\deg(v)\leq d$, it holds for all $i\in S_{v}$ that
$\widehat{\deg}_{\delta,i}(v)\leq\lambda\deg(v)+n^{\gamma}<\frac{n^{\alpha}}{\log n}.$
In particular, $\widehat{\deg}_{\delta,i}(v)<n^{\alpha}/\log(n)$ for at least half of the cascades. The correctness of Algorithm 3 follows. ∎
In the rest of this section, we focus on proving Lemma 4.1 through a discretization argument. Formally, we define the (finite) set of time indices
$\mathbb{T}_{n}:=\left\{\frac{k}{n^{5}}:k\in\operatorname{\mathbb{Z}}_{\geq 0}\right\}\cap[0,n^{2}].$
For $t\in\mathbb{T}_{n}$, let us also define the ground-truth discretized second derivative
@@
$dI_{\beta}(t):=\lambda\left(\operatorname{\mathsf{cut}}({\mathcal{I}}(t))-\operatorname{\mathsf{cut}}\left({\mathcal{I}}\left(t-\frac{1}{n^{5}}\right)\right)\right),$
as well as the empirical second derivative
$\widehat{dI_{\beta}}(t):=n^{\beta}\left(I\left[t,t+\frac{1}{n^{\beta}}\right]-I\left[t-\frac{1}{n^{\beta}},t\right]\right).$
Note in particular that $\widehat{dI_{\beta}}(T(v))=\widehat{\deg}_{\delta}(v)$ with $\delta=n^{-\beta}$. Our strategy is to show that for any fixed $t\in\mathbb{T}_{n}$, $\widehat{dI_{\beta,i}}(t)$ is close to $dI_{\beta,i}(t)$ in at least half of the cascade traces. We will then take a union bound over all $t\in\mathbb{T}_{n}$. Finally, Lemma 4.1 follows from showing that the discretization in $\mathbb{T}_{n}$ is fine enough so that $\widehat{\deg}_{\delta,i}(v)$ is close to $\widehat{dI_{\beta,i}}(t)$ and $dI_{\beta,i}(t)$ is close to $\lambda\deg(v)$ for some $t\in\mathbb{T}_{n}$ and all $i\in[K]$.
To execute this strategy, we start by showing that $\widehat{dI_{\beta}}(t)$ and $dI_{\beta}(t)$ are close for a fixed $t\geq 0$, with high probability. We also recall the definition of the event ${\mathcal{A}}$ from Definition 3.4, which we use in the lemma.
Lemma 4.2.
Let $G\in{\mathcal{G}}$, and let $\beta,\gamma$ satisfy $\beta>1/2$ and $\gamma>(1+\beta)/2$. For any $t\in\mathbb{T}_{n}$, it holds that
($\operatorname{\mathbb{P}}\left(\left\{\left|\widehat{dI_{\beta}}(t)-dI_{\beta}(t)\right|\geq n^{\gamma}\right\}\cap{\mathcal{A}}\right)\leq n^{-(\beta-1/2)+o(1)}.$
Since the proof is somewhat involved, we defer the details to Section 5.
Notice that the probability bound in Lemma 4.2 is not small enough to handle a union bound over the $O(n^{7})$ elements of $\mathbb{T}_{n}$. In the following result, we therefore consider multiple cascade traces to sufficiently reduce the probability of “bad” events.
Lemma 4.3.
Suppose that $\beta,\gamma$ satisfies $\beta>1/2$ and $\gamma>(1+\beta)/2$. Additionally assume
$K\geq\frac{15}{\beta-1/2}.$
Then with probability $1-o(1)$, there exists a set $S_{t}\subset[K]$ for each $t\in\mathbb{T}_{n}$ such that $|S_{t}|\geq\frac{K}{2}$ and
$\left|\widehat{dI_{\beta,i}}(t)-dI_{\beta,i}(t)\right|\leq n^{\gamma},\qquad i\in S_{t}.$
In words, it holds for all $t\in\mathbb{T}_{n}$ that in at least half of the cascade traces, $\widehat{dI_{\beta}}(t)$ is a good approximation for $dI_{\beta}(t)$. However, it need not be the case that the same cascade provides a good approximation for all $t$ simultaneously.
Proof of Lemma 4.3.
For $i\in[K]$, let ${\mathcal{A}}_{i}$ be the event ${\mathcal{A}}$ corresponding to the cascade ${\mathcal{I}}_{i}$. Observe that on the complement of the event described in Lemma 4.3, there must exist $t\in\mathbb{T}_{n}$ and a set $S\subset[K]$ with $|S|\geq K/2$ such that $|\widehat{dI_{\beta,i}}(t)-dI_{\beta,i}(t)|\geq n^{\gamma}$, for all $i\in S$. For a fixed $t\in\mathbb{T}_{n}$, we can bound this event as follows.
($\operatorname{\mathbb{P}}\left(\left\{\exists S\subset[K]:|S|\geq\frac{K}{2}\text{ and }\min_{i\in S}|\widehat{dI_{\beta,i}}(t)-dI_{\beta,i}(t)|\geq n^{\gamma}\right\}\cap\bigcap_{i\in[K]}{\mathcal{A}}_{i}\right)$
$\hskip 28.45274pt\leq\sum_{S\subset[K]:|S|\geq K/2}\operatorname{\mathbb{P}}\left(\left\{\min_{i\in S}|\widehat{dI_{\beta,i}}(t)-dI_{\beta,i}(t)|\geq n^{\gamma}\right\}\cap\bigcap_{i\in S}{\mathcal{A}}_{i}\right)$
=$\hskip 28.45274pt\stackrel{{\scriptstyle(a)}}{{=}}\sum_{S\subset[K]:|S|\geq K/2}\prod_{i\in S}\operatorname{\mathbb{P}}\left(\left\{|\widehat{dI_{\beta,i}}(t)-dI_{\beta,i}(t)|\geq n^{\gamma}\right\}\cap{\mathcal{A}}_{i}\right)$
$\hskip 28.45274pt\stackrel{{\scriptstyle(b)}}{{\leq}}\sum_{S\subset[K]:|S|\geq K/2}n^{-(\beta-1/2)K/2+o(1)}\stackrel{{\scriptstyle(c)}}{{=}}o\left(n^{-7}\right).$
Above, $(a)$ follows from the independence of the $K$ cascades, $(b)$ uses Lemma 4.2, and $(c)$ follows from our choice of $K$ and since the number of subsets of $[K]$ is $2^{K}=n^{o(1)}$. Taking a union bound over the $n^{7}$ elements of $\mathbb{T}_{n}$ shows that
($\operatorname{\mathbb{P}}\left(\left\{\exists t\in\mathbb{T},S\subset[K]:|S|\geq\frac{K}{2}\text{ and }\min_{i\in S}|\widehat{dI_{\beta,i}}(t)-dI_{\beta,i}(t)|\geq n^{\gamma}\right\}\cap\bigcap_{i\in[K]}{\mathcal{A}}_{i}\right)=o(1).$
Finally, the desired result follows since
$\operatorname{\mathbb{P}}\left(\left(\bigcap_{i\in[K]}{\mathcal{A}}_{i}\right)^{c}\right)\leq\sum_{i=1}^{K}\operatorname{\mathbb{P}}({\mathcal{A}}_{i}^{c})=o(1),$
where the final expression is due to Lemma 3.7. ∎
Next, we argue that $t\in\mathbb{T}_{n}$ can be replaced with $\{T(v)\}_{v\in V}$ in the statement of Lemma 4.3, which will prove Lemma 4.1. Towards this goal, we show that the discretization $\mathbb{T}_{n}$ is rich enough to capture the continuous-time dynamics of the cascade.
Lemma 4.4.
With probability $1-o(1)$, for each $v\in V$ there exists a unique $t\in\mathbb{T}_{n}$ such that $t-1/n^{5}\leq T(v)\leq t$.
@@
We first show that $\max_{v\in V}T(v)\leq n^{2}$. From the representation in Proposition 3.1, we can upper bound $\max_{v\in V}T(v)$ by $\sum_{e\in E(G)}F(e)$, so we focus on characterizing the latter quantity. In particular, we have that
$\operatorname{\mathbb{E}}\left[\sum_{e\in E(G)}F(e)\right]=\frac{E(G)}{\lambda}.$
We can upper bound the expectation by noting that $E(G)\leq dn+mn=n^{1+o(1)}$, which holds since there are at most $n$ vertices of degree at most $d$ and at most $m$ vertices of degree at most $n$ for $G\in{\mathcal{G}}(n,m,d,D)$. Markov’s inequality now implies
$\operatorname{\mathbb{P}}\left(\max_{v\in V}T(v)\geq n^{2}\right)\leq n^{-1+o(1)}.$
Hence there must exist some interval of the form $[t,t+1/n^{5}]$ which contains $T(v)$, for every $v\in V$.
We next show that in each interval of the form $[t,t+1/n^{5}]$, there exists at most one infection event. Condition on ${\mathcal{F}}_{t}$. If there are at least 2 infection events in $[t,t+1/n^{5}]$, then there must exist distinct $e_{1},e_{2}\in E(G)$ such that $F(e_{1}),F(e_{2})\leq n^{-5}$ (if not, at most one vertex is infected in the interval). The probability of this occurring is at most
(${|E(G)|\choose 2}\left(1-e^{-\lambda/n^{5}}\right)^{2}\leq n^{2+o(1)}\left(\frac{\lambda}{n^{5}}\right)^{2}=n^{-8+o(1)}.$
We conclude by taking a union bound over the $O(n^{7})$ elements of $\mathbb{T}_{n}$. ∎
We are now ready to prove the main result.
Proof of Lemma 4.1.
Assume that the events described in Lemmas 4.3 and 4.4 as well as ${\mathcal{A}}_{i}$ hold for all $i\in[K]$, which is the case with probability $1-o(1)$. Additionally, for a given $v\in V$, let $t\in\mathbb{T}_{n}$ satisfy $t-n^{-5}\leq T_{i}(v)\leq t$. Then we have that
$n^{-\beta}\left(\widehat{\deg}_{\delta,i}(v)-\widehat{dI_{\beta,i}}(t)\right)$ $=2I[T_{i}(v),t]-I\left[T_{i}(v)+\frac{1}{n^{\beta}},t+\frac{1}{n^{\beta}}\right]-I\left[T_{i}(v)-\frac{1}{n^{\beta}},t-\frac{1}{n^{\beta}}\right].$
On the event described in Lemma 4.4, no other vertex is infected in the interval $[t-n^{-5},t]$, hence $I[T_{i}(v),t]=1$. Furthermore, since the interval $[T(v)+n^{-\beta},t-n^{-\beta}]$ intersects at most two intervals of the form $[k/n^{5},(k+1)/n^{5}]$, the second term on the right hand side is at most 2 in absolute value. The same argument holds for the interval $[T_{i}(v)-n^{-\beta},t+n^{-\beta}]$. As a result, we have that
$\left|\widehat{\deg}_{\delta,i}(v)-\widehat{dI_{\beta,i}}(t)\right|\leq 6n^{\beta}.$
On the other hand, since $v$ is the only vertex which becomes infected in the interval $[t-n^{-5},t]$, we have that
$\operatorname{\mathsf{cut}}({\mathcal{I}}_{i}(t))=\operatorname{\mathsf{cut}}({\mathcal{I}}_{i}(T_{i}(v)))\text{ and }\operatorname{\mathsf{cut}}\left({\mathcal{I}}_{i}\left(t-\frac{1}{n^{5}}\right)\right)=\operatorname{\mathsf{cut}}({\mathcal{I}}(T_{i}(v)^{-})).$
As a result,
$dI_{\beta,i}(t)$ $=\lambda\left(\operatorname{\mathsf{cut}}({\mathcal{I}}_{i}(T_{i}(v)))-\operatorname{\mathsf{cut}}({\mathcal{I}}_{i}(T_{i}(v)^{-}))\right)=\lambda\left(\deg(v)-2|{\mathcal{N}}(v)\cap{\mathcal{I}}_{i}(T_{i}(v))|\right),$
where the second equality is due to (2.1).
On the event $\bigcap_{i\in[K]}{\mathcal{A}}_{i}$, it holds that $|{\mathcal{N}}(v)\cap{\mathcal{I}}_{i}(T_{i}(v))|\leq n^{1/2+o(1)}$. Putting everything together, it holds with probability $1-o(1)$ that
$\left|\widehat{\deg}_{\delta,i}(v)-\lambda\deg(v)\right|\leq\left|\widehat{dI_{\beta,i}}(t)-dI_{\beta,i}(t)\right|+6n^{\beta}+n^{1/2+o(1)}.$
In particular, for $i\in S_{t}$ we have that $|\widehat{\deg}_{i}(v)-\lambda\deg(v)|\leq 3n^{\gamma}$. As this holds for all $v\in V$, the desired result follows. ∎
5 Concentration of the second derivative: Proof of Lemma 4.2
The proof of Lemma 4.2 relies on the following two results. Both results use the event ${\mathcal{A}}$ (see Definition 3.4).
@@
Let $G\in{\mathcal{G}}$, let $\beta>1/2$ and assume that $\gamma>(1-\beta)/2$. Then for any $t\geq 0$, it holds that
($\operatorname{\mathbb{P}}\left(\left\{\left|I\left[t,t+\frac{1}{n^{\beta}}\right]-\frac{\lambda}{n^{\beta}}\operatorname{\mathsf{cut}}({\mathcal{I}}(t))\right|\geq n^{\gamma}\right\}\cap{\mathcal{A}}\right)\leq n^{-(\beta-1/2)+o(1)}.$
Lemma 5.2.
Let $\beta>1/2$ and $\gamma>\beta$. Then
($\operatorname{\mathbb{P}}\left(\left\{\left|\operatorname{\mathsf{cut}}\left({\mathcal{I}}\left(t-\frac{1}{n^{5}}\right)\right)-\operatorname{\mathsf{cut}}\left({\mathcal{I}}\left(t-\frac{1}{n^{\beta}}\right)\right)\right|\geq n^{\gamma}\right\}\cap{\mathcal{A}}\right)\leq n^{-(\beta-1/2)+o(1)}.$
Before proving the two lemmas, we show how they can be combined to prove the main result.
Proof of Lemma 4.2.
Let $\gamma>\beta$, and define the events
${\mathcal{E}}_{1}$ :=$:=\left\{\left|I\left[t,t+\frac{1}{n^{\beta}}\right]-\frac{\lambda}{n^{\beta}}\operatorname{\mathsf{cut}}({\mathcal{I}}(t))\right|\leq n^{\gamma-\beta}\right\}$
${\mathcal{E}}_{2}$ :=$:=\left\{\left|I\left[t-\frac{1}{n^{\beta}},t\right]-\frac{\lambda}{n^{\beta}}\operatorname{\mathsf{cut}}\left({\mathcal{I}}\left(t-\frac{1}{n^{\beta}}\right)\right)\right|\leq n^{\gamma-\beta}\right\}$
${\mathcal{E}}_{3}$ :=$:=\left\{\left|dI_{\beta}(t)-\lambda\left(\operatorname{\mathsf{cut}}({\mathcal{I}}(t))-\operatorname{\mathsf{cut}}\left({\mathcal{I}}\left(t-\frac{1}{n^{\beta}}\right)\right)\right)\right|\leq n^{\gamma}\right\}.$
On ${\mathcal{E}}_{1}\cap{\mathcal{E}}_{2}$ it holds that
$\left|\widehat{dI_{\beta}}(t)-\lambda\left(\operatorname{\mathsf{cut}}({\mathcal{I}}(t))-\operatorname{\mathsf{cut}}\left({\mathcal{I}}\left(t-\frac{1}{n^{\beta}}\right)\right)\right)\right|\leq 2n^{\gamma}.$
Additionally if ${\mathcal{E}}_{3}$ holds, we have by the triangle inequality that $|\widehat{dI_{\beta}}(t)-dI_{\beta}(t)|\leq 3n^{\gamma}=n^{\gamma+o(1)}$. It therefore suffices to study the probability of ${\mathcal{E}}_{1}\cap{\mathcal{E}}_{2}\cap{\mathcal{E}}_{3}$. By Lemma 5.1, it holds for $\gamma>(1+\beta)/2$ that
$\operatorname{\mathbb{P}}\left(({\mathcal{E}}_{1}\cap{\mathcal{E}}_{2})^{c}\cap{\mathcal{A}}\right)\leq\operatorname{\mathbb{P}}({\mathcal{E}}_{1}^{c}\cap{\mathcal{A}})+\operatorname{\mathbb{P}}({\mathcal{E}}_{2}^{c}\cap{\mathcal{A}})\leq n^{-(\beta-1/2)+o(1)}.$ (5.1)
It remains to study ${\mathcal{E}}_{3}$. Lemma 5.2 as well as the definition of $dI_{\beta}(t)$ implies that
$\operatorname{\mathbb{P}}({\mathcal{E}}_{3}^{c}\cap{\mathcal{A}})=\operatorname{\mathbb{P}}\left(\left\{\left|\operatorname{\mathsf{cut}}\left({\mathcal{I}}\left(t-\frac{1}{n^{5}}\right)\right)-\operatorname{\mathsf{cut}}\left({\mathcal{I}}\left(t-\frac{1}{n^{\beta}}\right)\right)\right|\geq n^{\gamma}\right\}\cap{\mathcal{A}}\right)\leq n^{-(\beta-1/2)+o(1)}.$ (5.2)
Together, the bounds (5.1) and (5.2) prove the lemma. ∎
The rest of this section is devoted to the proofs of the two lemmas. The proof of Lemma 5.1 is in Section 5.1, and the proof of Lemma 5.2 is in Section 5.2.
5.1 Approximating the cut: Proof of Lemma 5.1
We start by recalling and defining some notation. For $0\leq a\leq b$, recall that ${\mathcal{I}}[a,b]$ denotes the set of vertices infected in the interval $[a,b]$. In particular, $I[a,b]=|{\mathcal{I}}[a,b]|$. For $t\geq 0$, we also define the infection boundary
$\partial{\mathcal{I}}(t):=\left\{v\in V\setminus{\mathcal{I}}(t):{\mathcal{N}}(v)\cap{\mathcal{I}}(t)\neq\emptyset\right\}.$
In words, $\partial{\mathcal{I}}(t)$ is the set of vertices which are not yet infected at time $t$, but have at least one infected neighbor. Conditioned on ${\mathcal{F}}_{t}$, we partition ${\mathcal{I}}[t,t+n^{-\beta}]={\mathcal{J}}_{1}\cup{\mathcal{J}}_{2}$, where
${\mathcal{J}}_{1}$ :=$:=\left\{v\in\partial{\mathcal{I}}(t):\min_{u\in{\mathcal{N}}(v)\cap{\mathcal{I}}[0,t]}F(u,v)\leq\frac{1}{n^{\beta}}\right\}$
${\mathcal{J}}_{2}$ :=$:=\left\{v\in{\mathcal{I}}\left[t,t+\frac{1}{n^{\beta}}\right]:P\in{\mathcal{P}}(v,t)\text{ and }\mathsf{weight}(P)\leq\frac{1}{n^{\beta}}\Rightarrow|P|\geq 2\right\}={\mathcal{I}}\left[t,t+\frac{1}{n^{\beta}}\right]\setminus{\mathcal{J}}_{1}.$
We proceed by studying the sizes of ${\mathcal{J}}_{1}$ and ${\mathcal{J}}_{2}$ separately. Our first result characterizes $|{\mathcal{J}}_{1}|$, essentially through an application of Bernstein’s inequality.
Lemma 5.3.
Let $G$ be a graph on $n$ vertices, let $\beta>1/2$ and assume that $\gamma>(1-\beta)/2$. Recall the event ${\mathcal{A}}$ from Definition 3.4. Then for any $t\geq 0$, it holds that
($\operatorname{\mathbb{P}}\left(\left\{\left||{\mathcal{J}}_{1}|-\frac{\lambda}{n^{\beta}}\operatorname{\mathsf{cut}}({\mathcal{I}}(t))\right|\geq n^{\gamma}\right\}\cap{\mathcal{A}}\right)=o\left(\frac{1}{n}\right).$